OpenStudy (anonymous):

what is the diference between -1/10 and -8/10

OpenStudy (calculusxy):

\[-\frac{ 1 }{ 10 } - (-\frac{ 8 }{ 10 }) = -\frac{ 1 }{ 10 } + \frac{ 8 }{ 10 }\]

OpenStudy (calculusxy):

Now subtract and take the sign of the larger absolute value.